OpenStudy (icalibear):

Simplify the expression (-4)^-6(-4)^7?

OpenStudy (icalibear):

Yeah

OpenStudy (anonymous):

Use: \[\large x^a \cdot x^b = x^{a+b}\] Can you use this ??

OpenStudy (icalibear):

is the answer 4^13 ??

OpenStudy (anonymous):

Here x is -4 a = -6 and b = 7. No.. Try one more time...

OpenStudy (icalibear):

so its -4

OpenStudy (icalibear):

or -4^1

OpenStudy (anonymous):

\[(-4)^1\] It looks good to me.. But what is actually \((-4)^1\) ?

OpenStudy (icalibear):

-4

OpenStudy (anonymous):

Yep..

OpenStudy (icalibear):

how did i not get that the first time ._. lol thank you

OpenStudy (anonymous):

Because you are adding 6 and 7 neglecting the negative sign in front of 6. Like: 6 + 7 = 13 But in actual: -6 + 7 = 1..

OpenStudy (phi):

how did i not get that the first time ._. by ignoring detestable minus signs. They are ugly, so you naturally ignore them, hoping they will go away..

OpenStudy (icalibear):

yes ._.

OpenStudy (anonymous):

Did you get everything ?? @icalibear